# Photoluminescence of Femtosecond Laser-irradiated Silicon Carbide

### v1

@@ III Irradiation of pristine HPSI SiC
We then perform PL spectroscopy of the bright spots created by irradiation with single laser pulses. For all photo-luminescent spots, we obtain typical spectra as displayed in Fig. 2b). This very broad PL spectrum without a discernible zero-phonon line has been identified as the typical PL signature of at room temperature Carter et al. (2015). The measured PL intensities and the fact that second order photon correlation measurements on the bright spots show no anti-bunching point to ensembles and are similar to previous work Castelletto et al. (2018).
In FIG. 1 c), we analyze the spatial dimension of the laser irradiated spots on our PL map. To this end, we fit the PL spots using a 2D Gaussian function to obtain their diameter (1/) and plot it in dependence of the laser pulse energy. We note that for higher pulse energies the shape of the luminescent spots deviates from Gaussian, however, the Gaussian fit is considered a good estimate to extract the diameter. PL spots grow in size up to a pulse energy of 1590 nJ, subsequently reaching a plateau. This finding is in accordance with our AFM measurements that show a similar trend. For comparison with AFM measurements, a linear estimation gives a growth rate of the PL spots with laser pulse energy of 1.00 nm/nJ (0.41 nm/nJ growth rate with AFM measurements).
@@
We measure PL saturation curves for each laser irradiated spot and average over the measurements on the spots irradiated with the same pulse energy. Figure 3 a) exemplary shows the saturation measurement for spots irradiated with 230 nJ. To correct the data for background, we perform a saturation measurement at a spot 10 m downward from the row written at 60 nJ. We correct the data points by subtracting the background. We note that the PL background also showed a saturation behavior. This might indicate that native point defects contribute to the background as reported previously Castelletto et al. (2020). Figure 3. b) shows the saturation count rate as a function of the writing pulse energy. Saturation count rates increase linearly with laser pulse energy indicating a linear increase of luminescent defects with pulse energy assuming a constant emission rate for each created defect. On the other hand, we will below discuss changes on excited state lifetimes that indicate non-radiative processes and thus changes in the emission rate per defect. the decreasing lifetimes pointing out to non-radiative decaying pathways for make it complicated to estimate the evolution on the number of with laser pulse energy. We also have to take into account the scattering of light in the highly damaged regions influencing the collected photoluminescence.
@@
In addition modifying the surface profile (FIG. 1), increasing the pulse energy also changes the crystalline properties of SiC in the vicinity of the spots Zhou et al. (2023); Liu et al. (2020); Castelletto et al. (2018). When aiming to create , due to their lower formation energy, carbon vacancies () are also created nearby, which are not optically active. They provide non-radiative decay paths for therefore quenching their PL Steeds et al. (2002); Kraus et al. (2017); Chen et al. (2019). FIG.3. d) summarizes the lifetimes extracted from PL measurements (> 850 nm). Again, we obtain by averaging measurements from all spots irradiated with the same pulse energy. For spots irradiated with 230 nJ pulses, we find which matches obtained for created using electron irradiation Hain et al. (2014). linearly decreases with pulse energy, indicating quenching by and also potential interaction with extended defects due to laser irradiation. The decrease in renders it challenging to estimate number of luminescent defects it might also indicate a change in the type of defect being created. We also have to take into account the scattering of light in highly damaged regions influences PL collection. In contrast, the saturation power as displayed in FIG.3 c) did not change with pulse energy.
FIG. 2. c) shows Raman spectra taken using a commercial Raman spectrometer using a 532 nm laser light on the pristine sample from an unprocessed area and from spots written at 60 nJ, 230 nJ and 1850 nJ. Compared to the reference Raman spectra from the unprocessed 4H-SiC, we observe on spectra taken from dimples made with single pulse at 60 nJ slight drop in peak intensities most notably on the TO() and LO() peaks. At 230 nJ and 1850 nJ, the decrease in peak intensities is even more significant being maximal for the 1850 nJ. While this Raman spectrum shows no sign of amorphous material even for 1850 nJ, the strong decrease in Raman peak intensity has been associated with significant crystal damage in previous work Castelletto et al. (2018).
IV Irradiation of HPSI with graphene layer
The second sample with an epitaxial graphene layer grown on top of a HPSI 4H-SiC substrate was processed using the same method described in the previous section, except using for this sample a repetition rate of 100 kHz. Figure 4 a) shows a white-light microscope image of a laser processed pattern, having the same layout as presented in FIG. 1 a). In the single pulse region, we reduced the laser pulse energy to range from 60 nJ (bottom) to 230 nJ (top) or alternatively from 230 nJ to 520 nJ (pattern not shown). For this sample, we also created patterns for which we lifted the sample stage by 5 m or 10 m respectively, to move the laser focus into the bulk of our SiC sample. FIG. 4. b) shows a room temperature confocal PL map of the pattern from FIG. 4. a). The background PL is more pronounced compared to FIG. 2 a). Thus, the annealing and growth of the graphene layer did not aid in removing background PL in our case. We also note as a general finding that patterns deeper in the bulk show less PL than the patterns
To obtain more information on the defects created by the laser writing process, the sample with graphene was cooled down to 4.3K in a closed-cycle cryostat (CryoVac). In this cryogenic confocal setup, we excite PL using a tunable titanium sapphire laser (Sirah) operating at 737.19 nm focused on the sample with a microscope objective (Olympus MPLN 100x, NA = 0.9). PL light is collected using the same objective and reflected laser light is separated from the PL with a 740 nm dichroic mirror. Transmitted PL light is further filtered with a longpass filter and coupled to a single mode fiber that can either be connected to an APD (Excelitas SPCM-AQRH) or a spectrometer (Horiba iHR550) equipped with a 600 lines per mm grating and a Peltier cooled CCD camera.
We record various PL spectra at 4.3K. While our room temperature spectra as well as the lifetime measurements supported the localized creation of , we do not observe the ZPL emission of at cryogenic temperature and our findings do thus not confirm an efficient, localized creation of . This finding is in accordance with Castelletto et al. (2020). Ref. Castelletto et al. (2020), reports in laser written spots but only a weak enhancement compared to their non-irradiated substrate. We note that potential in our substrate have been most probably eliminated due to the annealing in the graphene growth. Figure 4 c) compares cryogenic spectra recorded on SiC irradiated with a single 520 nJ pulse to non-irradiated SiC next to the pattern. Both spectra shows peaks associated to the TS center at 769 nm, 812 nm and 813 nm (TS1, TS2 and TS3 Rühl et al. (2018)). To investigate differences between defect creation at the surface and in the bulk, we investigate the pattern created by lifting the sample 10 m. We find two distinct peaks when scanning the depth of the excitation laser focus (Figure 6 inset). We take spectra on one of the markers irradiated with multiple pulses as PL from spots irradiated with single pulses was too weak. Spectra taken in the depths corresponding to the two PL maxima reveal again the signature of the TS and additional broadband PL Figure 6. Raman lines in both spectra have similar intensity and width, hinting at similar levels of crystal damage. So we cannot conclude on any differences in defect creation when the laser focus is irradiating the SiC surface or the bulk. The TS defect is known to survive high annealing temperature, however, its microscopic origin remains unclear. We thus cannot use TS defects to learn more about the laser writing process. The absence of could indicate very high temperatures in the processed volume that anneal out . Additionally, have to be in their negatively-charged state to display the desired PL. So changes in the charge state due to the presence of other defects might also de-active PL.
For the sample with the graphene layer, we also observe a strong increase in saturation power observed at room temperature. Figure 5. a)-b) exemplary compare saturation measurements for the PL created by the 230 nJ and 520 nJ pulse. Saturation count rate in the pristine(graphene) sample were measured to be 139(204) kcounts/s and 351(259) kcounts/s, for 230 and 520 nJ respectively. Strikingly, the associated saturation powers are 1.38 mW and 5.21 mW in the pristine sample and 25.6 mW and 18.5 mW in the sample with graphene. Neither the strong increase in saturation power nor the loss in PL can be connected to the absorption of the graphene layer which is only 2.3% Rühl et al. (2020).
@@ IV Irradiation of HPSI with graphene layer
We investigated the effect of epitaxial graphene on HPSI SiC on the creation of luminescent defects under femtosecond laser irradiation. In accordance with previous work, we do not observe efficient creation of centers, while spots with intense PL seem to be connected with strong surface structures (dimples) in a reference sample without graphene. Despite strong surface modification, we do not see signature of amorphous materials. To optimize the process, 4H-SiC with a more favorable crystal orientation and doping could be investigated as well as shorter laser wavelength and stronger focusing to enhance non-linear effects in comparison to sample heating.

### v2

@@ IV Irradiation of pristine HPSI SiC
We then perform PL spectroscopy of the bright spots created by irradiation with single laser pulses. For all photo-luminescent spots, we obtain typical room temperature spectra as displayed in figure 3 c). This very broad PL spectrum without a discernible zero-phonon line has been identified as the typical PL signature of at room temperature Carter et al. (2015). The measured PL intensities and the fact that second order photon correlation measurements on the bright spots show no anti-bunching point defect ensembles. Room temperature spectra are similar to previous work Castelletto et al. (2018).
@@
We measure PL saturation curves for each laser irradiated spot and average over the measurements on the spots irradiated with the same pulse energy. Figure 5 a) exemplary shows the saturation measurement for spots irradiated with 230 nJ. To correct the data for background, we perform a saturation measurement at a spot 10 m downward from the row written at 60 nJ. We correct the data points by subtracting the background. We note that the PL background also showed a saturation behavior. This might indicate that native point defects contribute to the background as reported previously Castelletto et al. (2020). Figure 5 b) shows the saturation count rate as a function of the writing pulse energy. Saturation count rates increase linearly with laser pulse energy indicating a linear increase of luminescent defects with pulse energy assuming a constant emission rate for each created defect. On the other hand, we will below discuss changes on excited state lifetimes that indicate non-radiative processes and thus changes in the emission rates.
@@
In addition, modifying the surface profile (Figure 4), especially for increased pulse energy also changes the crystalline properties of SiC in the vicinity of the spots Zhou et al. (2023); Liu et al. (2020); Castelletto et al. (2018). When aiming to create , carbon vacancies () are also created nearby due to their low formation energy. Carbon vacancies are not optically active. They provide non-radiative decay paths for therefore quenching their PL Steeds et al. (2002); Kraus et al. (2017); Chen et al. (2019). Figure5. d) summarizes the lifetimes extracted from PL measurements (> 850 nm). Again, we obtain by averaging measurements from all spots irradiated with the same pulse energy. For spots irradiated with 230 nJ pulses, we find ns which matches ns obtained for created using electron irradiation Hain et al. (2014) again hinting at the creation of . linearly decreases with pulse energy, indicating quenching by and also potential interaction with extended defects due to laser irradiation. In contrast, the saturation power as displayed in figure 5 c) did not change with pulse energy.
To obtain more information on the defects created by the laser writing process, both samples (graphene and pristine) and a reference HPSI 4H-SiC sample implanted with 350 keV protons at a dose of were cooled down to 4.3K in the CLSM 1 setup described in the previous section.
Figure 6 a) compares the cryogenic spectra of three 4H-SiC samples: ion-implanted reference sample, laser written pristine sample, and laser written sample with a graphene layer. The reference sample exhibits peaks associated with silicon vacancies (), including V1’ at 858 nm, V1 at 861 nm, and V2 at 916 nm, in agreement with literature.
While our room-temperature spectra and lifetime measurements strongly suggested the localized formation of (see figure 3 d) and figure 5 d)), we do not observe ZPL emission from at cryogenic temperatures in both laser-written samples. This suggests that efficient, localized creation of is not feasible under our experimental conditions and using HPSI material. Our findings align with those reported in Castelletto et al. (2020), where centers were detected in laser-written regions but only with a weak enhancement relative to the non-irradiated substrate. Moreover, as summarized in table 1 previous laser writing work performed at the same 1030 nm wavelength did not show ZPL related to , only PL creation and typical room temperature spectra were observed. centers occurring before laser writing in our substrates were likely eliminated due to the high-temperature annealing for the pristine sample and the graphene growth for the other sample.
The PL spectrum of the sample with graphene exhibits peaks corresponding to the TS center at 769 nm, 812 nm, and 813 nm (TS1, TS2, and TS3, respectively) Rühl et al. (2018). To determine whether TS centers form during laser writing or graphene growth, we analyzed spectra from a laser-written spot (520 nJ) and an unprocessed region of the substrate, as shown in figure 6 b). The absence of significant spectral differences suggests that TS centers originate from the graphene growth process and not from laser irradiation. Additionally, the Raman spectra of both regions exhibits comparable intensity and linewidth, indicating similar levels of crystal damage. The TS defect is known to withstand high annealing temperatures, but its microscopic origin remains uncertain and TS centers cannot serve as indicators for studying the laser-writing process. The absence of in laser-written regions may suggest that high temperatures in the processed volume lead to their annealing. Moreover, for centers to exhibit PL, they must be in their negatively charged state. Charge-state alterations caused by the presence of other defects may therefore suppress PL emission.
Moreover, we investigated the near-infrared (NIR) PL using a dedicated setup optimized for detection above 1000 nm. Within this wavelength range falls the PL of divancancy defect in SiC. The samples were cooled down to 5 K in an attoDRY800 cryostat, and different excitation wavelengths were employed: 730 nm and 915 nm from a Sirah Ti:Sa Matisse CR laser, and 810 nm from a Toptica diode (EYP-RWL-0808-00800-4000 BFW09-0000). The excitation beams were focused onto the samples and PL was collected using a Zeiss 100×, NA 0.9 objective. Reflected laser light was removed from the collected PL using dichroic mirrors and a set of long-pass filters before being sent to a SpectraPro HRS 750 spectrometer equipped with a NIRvana HS camera.
The sample with a graphene layer was investigated under 810 nm excitation, however, no PL peaks associated with the divacancy were observed. In contrast, the pristine sample was studied under 915 nm excitation with a 1000 nm long-pass filter. Multiple spectra were recorded both from unprocessed bulk regions (reference) and from laser written spots at 1850 nJ, taken close to and far from markers, example spectra are shown in figure 7.
In the reference spectra from an unprocessed region, PL peaks characteristic of the divacancy are visible Koehl et al. (2011); Magnusson et al. (2018): PL1 (1132 nm), PL2 (1130 nm), and PL4 (1078 nm). These centers were either present natively in the crystal or formed during the annealing step at 800°C. At the laser written spots, the same spectral features can be identified, however, they appear broadened compared to the reference. This broadening is attributed to lattice damage induced by the laser writing process, and is most pronounced around the alignment markers, where the crystal damage is most severe. The absence of any PL intensity enhancement and the spectral broadening of the divacancy-related peaks, suggests that the laser writing process did not generate additional divacancies, but instead degraded or partially destroyed the ones already present before laser processing.
@@
We investigated the effect of epitaxial graphene on HPSI SiC on the creation of luminescent defects under femtosecond laser irradiation. We employ a commercial HPSI substrate and an industrial grade laser system to test broader applicability of the method. We observed a much lower intensity threshold to create PL defects on the sample with graphene compared to the pristine sample, suggesting graphene alters the local absorption and energy deposition conditions on the 4H-SiC surface during laser writing. In accordance with previous work, we do not observe efficient creation of or divacancies centers, while spots with intense PL seem to be connected with strong surface structures (dimples) in a reference sample without graphene. Despite strong surface modification, we do not see signature of amorphous materials. To optimize the process, 4H-SiC substrate with a 4H-SiC epitaxial layer with a more favorable doping could be investigated as well as shorter laser wavelength and stronger focusing to enhance non-linear effects in comparison to sample heating.